### v1

# DiffuseDef: Improved Robustness to Adversarial Attacks

Abstract
Pretrained language models have significantly advanced performance across various natural language processing tasks. However, adversarial attacks continue to pose a critical challenge to system built using these models, as they can be exploited with carefully crafted adversarial texts. Inspired by the ability of diffusion models to predict and reduce noise in computer vision, we propose a novel and flexible adversarial defense method for language classification tasks, DiffuseDef111https://github.com/Nickeilf/DiffuseDef, which incorporates a diffusion layer as a denoiser between the encoder and the classifier. During inference, the adversarial hidden state is first combined with sampled noise, then denoised iteratively and finally ensembled to produce a robust text representation. By integrating adversarial training, denoising, and ensembling techniques, we show that DiffuseDef improves over different existing adversarial defense methods and achieves state-of-the-art performance against common adversarial attacks.
DiffuseDef: Improved Robustness to Adversarial Attacks
Zhenhao Li, Marek Rei, Lucia Specia Language and Multimodal AI (LAMA) Lab, Imperial College London {zhenhao.li18, marek.rei, l.specia}@imperial.ac.uk
1 Introduction
Pretrained language models (PLM) have significantly advanced the performance of various natural language processing (NLP) tasks. Despite such improvements, current NLP systems remain susceptible to adversarial attacks where carefully crafted text perturbations can lead to incorrect model outputs (Alzantot et al., 2018; Jin et al., 2020; Li et al., 2020). In order to improve robustness to adversarial attacks, various defense methods have been proposed, such as adversarial training (Zhu et al., 2020; Si et al., 2021; Zhou et al., 2021; Xi et al., 2022), text denoising (Nguyen Minh and Luu, 2022; Wang et al., 2023), ensembling (Zhou et al., 2021; Zeng et al., 2023; Li et al., 2023), etc. However, existing defense methods either assume the test-time perturbation/attack set is similar to that used in training (Li et al., 2021), or are limited to specific architectures (Xi et al., 2022), or at inference time require large computational cost, thereby limiting their practical applicability.
Diffusion models are commonly used in computer vision (CV) to generate high-quality images by predicting and removing noise from a sampled noisy image. Therefore, they can be adopted to remove noise from adversarial images and thus improve robustness to attacks (Nie et al., 2022). However, in NLP very limited research has investigated adversarial defense with diffusion models due to the discrete and contextual nature of text data. Li et al. (2023) adopt the idea of iterative denoising and reconstruct adversarial texts from masked texts, while Yuan et al. (2024) use a diffusion model as a classifier and perform reverse diffusion steps on the label vector, conditioning on the input text. Inspired by the general noise prediction and reduction capability of diffusion models, we propose DiffuseDef, a novel adversarial defense method which employs diffusion training to denoise hidden representations of adversarial texts. Unlike Li et al. (2023) and Yuan et al. (2024) which apply diffusion on texts or labels, DiffuseDef directly removes noise to the hidden states, providing a more effective and robust text representation to defend against adversarial texts. Compared to diffusion-based defense in CV (Nie et al., 2022), DiffuseDef further enhances robustness with ensembling and improves efficiency with fewer diffusion steps.
DiffuseDef combines adversarial training with diffusion training, where the diffusion layer is trained to predict randomly sampled noise at a given timestep. During inference, the diffusion layer serves as a denoiser, iteratively removing noise from adversarial hidden states to yield a robust hidden representation. Moreover, we adopt the ensembling strategy by first adding random noise to text hidden states to create multiple variants then denoising them via the diffusion layer. The model output is made by averaging all denoised hidden states. Since ensembling happens solely at the diffusion layer, DiffuseDef is more efficient than traditional ensembling-based methods (Ye et al., 2020; Zeng et al., 2023), which require a full forward pass through all model parameters.
Through systematic experimentation, we demonstrate that DiffuseDef outperforms strong defense methods and is able to defend against multiple types of adversarial attacks, while preserving performance on clean texts. Our analysis also reveals that the ensembling diffused representations provides a stronger defense against finding vulnerable words to attack and can reduce the distance in latent space between adversarial texts and their clean text counterpart.
Our contributions can be summarised as follows:
We propose DiffuseDef, a novel and flexible adversarial defense method that can be added on top of any existing adversarial defense methods to further improve robustness to adversarial attacks.
DiffuseDef outperforms existing adversarial methods and achieves state-of-the-art performance against prevalent adversarial attacks.
Through extensive analysis, we demonstrate the effectiveness of the ensembling diffused representation and the efficiency of DiffuseDef compared to existing ensembling-based methods.
2 Related Work
2.1 Textual Adversarial Attacks
Textual adversarial attacks focus on constructing adversarial examples from an original text that maximise the likelihood of incorrect predictions by a neural network. These attacks require adversarial examples to be perceptually similar to the original text, which is typically achieved by introducing subtle perturbations to the original text, such as character swapping (Gao et al., 2018; Ebrahimi et al., 2018), synonym-substitutions (Ren et al., 2019; Yoo and Qi, 2021), and paraphrasing (Gan and Ng, 2019; Huang and Chang, 2021). Taking the text classification task as an example, given a classifier that maps an input sequence of words to its designated label , the goal of the attack model is to construct an adversarial example to fool the classifier, where is a subtle adversarial perturbation constrained by . The adversarial example is considered a successful attack if it leads to an incorrect prediction . The attacker can iteratively generate multiple adversarial examples and query the classifier to obtain a successful attack, whereas the classifier must consistently return the correct prediction within a specified number of query attempts to be considered robust.
Common textual adversarial attack methods adopt a two-stage process to construct effective adversarial examples: word importance ranking and word substitution. In the first stage, words or subwords are ranked based on their influence on the model’s prediction. This is measured by leveraging either gradient information (Liu et al., 2022) or changes in prediction probabilities when words are removed (Jin et al., 2020) or masked (Ren et al., 2019; Li et al., 2020). In the second stage, candidate words are substituted with synonyms (Zang et al., 2020), perturbed variants (Gao et al., 2018), or outputs from masked language models (Garg and Ramakrishnan, 2020; Li et al., 2020). The substitution process is guided by various constraints to ensure the adversarial example remains natural and semantically equivalent to the original text. Common constraints include thresholding the similarity between the replacement word embedding and the substituted word embedding, or ensuring the semantic similarity between sentence vectors modeled from Universal Sentence Encoder (Cer et al., 2018). Despite these constraints, current textual adversarial attacks still pose significant challenges to NLP models (Liu et al., 2022; Xu et al., 2021; Yuan et al., 2023), highlighting the necessity for defense methods for better adversarial robustness.
2.2 Adversarial Defense Methods
To mitigate the performance degradation caused by adversarial attacks, various adversarial defense methods have been developed. They can be grouped into three categories: training-based, ensembling-based, and denoising-based methods. Adversarial training improves the robustness of the model to adversarial examples through strategies like data augmentation (Si et al., 2021) and adversarial regularisation (Madry et al., 2018; Zhu et al., 2020; Wang et al., 2021; Xi et al., 2022; Gao et al., 2023). However, adversarial training methods are limited as they assumes similar train-test adversarial examples, and thus tend to overfit to specific types of adversarial attacks. Ensembling-based methods generate multiple variants of the input text at inference time and ensemble model predictions over all the variants (Ye et al., 2020; Zhou et al., 2021; Zeng et al., 2023; Li et al., 2023), but they can be inefficient given that model predictions are needed on every ensemble, increasing the inference time with the number of ensembles. More recently, denoising-based methods have been proposed to improve adversarial robustness by mapping the vector representation of the adversarial text to another point in the latent space that is close to the clean text (Nguyen Minh and Luu, 2022; Wang et al., 2023; Moon et al., 2023; Yuan et al., 2024). The denoised representation makes it more difficult to find vulnerable words to attack, thus improving adversarial robustness (Wang et al., 2023). Nevertheless, denoising might lead to very different representations of clean text and adversarial text, therefore changing the semantic meanings.
The proposed DiffuseDef builds on these three approaches and can use any adversarially trained classifier as the base, applying denoising via a diffusion layer, and ensembling the diffused representations with a small number of ensembles. Using a diffusion layer as a denoiser addresses the overfitting problem from adversarial training and mitigates the efficiency problem by performing ensembling only at the diffusion layer. By averaging denoised hidden states across all ensembles, DiffuseDef also addresses the issue stemming from denoising, maintaining good performance on clean texts.
3 DiffuseDef
3.1 Training
The proposed diffusion defense model consists of a pretrained encoder for feature extraction, a transformer-based diffusion layer for noise prediction and reduction, and a classifier layer for output generation. The training process is split into two stages: adversarial training and diffusion training (Figure 1). The adversarial training stage employs any neural network-based adversarial training methods like FreeLB++ (Li et al., 2021) and RSMI (Moon et al., 2023), which optimise the encoder and classifier for robustness by perturbing the latent representation of the text input.
In the diffusion training stage, only the diffusion layer is trained to predict random noise added to the clean text hidden state at different timesteps, enabling it to denoise the adversarial hidden state at inference time. The pretrained encoder, however is frozen during this stage. Since the pretrained encoder is only used for feature extraction, the diffusion training method is compatible with any neural network-based adversarial training method.
Given an input sequence of tokens , the pretrained encoder extracts the hidden state . A random Gaussian noise is sampled to perturb hidden state . Sohl-Dickstein et al. (2015) define the forward diffusion process as a Markov Chain where at each timestep a Gaussian noise is sampled and added to the previous latent feature: , where , is the noisy hidden state at step and is a pre-calculated variance schedule changing with . As shown by Ho et al. (2020), this equation can be reformulated to calculate directly from by defining and , thus
At each training step, a random forward diffusion timestep is sampled from a uniform distribution. Therefore, the noisy hidden state is created from , , and . The diffusion layer consists of a time embedding and a transformer layer. The time embedding receives the diffusion timestep as input and produces an embedding , which is added to as input for the transformer layer. Finally, the transformer layer outputs the predicted noise , and mean square error is used to compute the loss between predicted noise and actual sampled noise .
3.2 Inference
Leveraging the diffusion layer’s ability to predict noise at a given timestep , we utilise it as a denoiser during inference by iteratively performing the reverse diffusion steps, which sample from to produce the denoised hidden state
where is the predicted noise from diffusion layer and . The denoised hidden state can thus be computed with
where .
Inference in DiffuseDef combines a one-step noising, a multi-step denoising, and an ensembling step. After the pretrained encoder extracts its hidden state , a set of Gaussian noise vectors are sampled to perform a single forward diffusion step. These noise vectors are then added to the hidden state following equation 1, resulting in a set of noisy hidden states , where denotes the number of denoising steps. The noisy hidden states are subsequently denoised through reverse diffusion steps, where noise is predicted by the diffusion layer and subtracted from the previous noisy hidden states. Unlike Ho et al. (2020) where the reverse diffusion step starts with pure noise sampled from standard normal distribution, we assume the noisy hidden state is already an intermediate state in the reverse diffusion steps. This allows us to use a smaller number of than the training timestep to prevent the denoised hidden states from diverging substantially from the initial hidden state . This sequence of denoising steps creates the final denoised hidden states , which are averaged and used by the classifier to output the final predicted label. This process is summarised in Algorithm 1.
4 Experiments
Datasets
We focus on two common NLP tasks in our experiments: topic classification and natural language inference (NLI). In the text classification task, we compare our method with other defense algorithms on two standard datasets for adversarial defense: AG News (Zhang et al., 2015a) and IMDB (Maas et al., 2011a) datasets. In the NLI task, we perform an ablation analysis with the Question-answering NLI (QNLI) dataset (Wang et al., 2018). We randomly split AGNews, IMDB, and QNLI datasets into train, validation, and test splits.
Evaluation
Following previous work on adversarial defense, we use three benchmarking attack methods to evaluate the robustness of DiffuseDef: TextFooler (TF) (Jin et al., 2020), TextBugger (TB) (Li et al., 2019), and Bert-Attack (BA) (Li et al., 2020). The three attack methods create adversarial attacks in different granularities: character-level perturbation (TextBugger), word substitution (TextFooler), and subword substitution (BertAttack). Regarding evaluation metrics, we measure the clean accuracy (Clean%) on the test set, the accuracy under attack (AUA%), and the number of adversarial queries (#Query) needed for a successful attack. Higher scores on the three metrics denote a better robustness performance of a defense method. The accuracy on clean data is measured across the entire test set. The accuracy under attack and number of queries, due to the lengthy attacking process, is measured on a randomly sampled subset of 1000 examples from the test set. We use the TextAttack library as the adversarial evaluation framework. To ensure a fair comparison and high-quality adversarial examples, we follow the same evaluation constraints as in Li et al. (2021). The evaluation metrics are averaged based on experiments run with 5 random seeds.
4.1 Comparison to SOTA
We compare our proposed method with state-of-the-art adversarial defense approaches, trained using both BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) as backbones: Fine-tune: Fine-tuning pretrained models on downstream task with no defense method applied222”Fine-tune” is a baseline approach used to illustrate the effect of adversarial attacks.. InfoBERT (Wang et al., 2021): Applying mutual-information-based regularizers during fine-tuning of pretrained models to improve robustness. FreeLB++ (Li et al., 2021): An adversarial training method improving on FreeLB(Zhu et al., 2020), which adds adversarial perturbations to word embedding during fine-tuning. EarlyRobust333We only run EarlyRobust with BERT as its implementation with RoBERTa has not been released. (Xi et al., 2022): Extracting early-bird subnetworks and pruning pretrained models for efficient adversarial training. RSMI (Moon et al., 2023): A two-stage training method that combines randomised smoothing and masked inference to improve adversarial robustness.
4.2 Implementation and Settings
We train two DiffuseDef variants using FreeLB++ and RSMI models as base models considering their robust adversarial defense capabilities. In the diffusion layer, only one transformer encoder layer (Vaswani et al., 2017) is used. The maximum noising timestep during training is set to 30 for AGNews and QNLI datasets, and 10 for IMDB dataset, while at inference time, we only apply 5 denoising steps for . We follow (Ho et al., 2020) to use a linear schedule from to . The diffusion layer is trained for 100 epochs, with the base classifier parameters frozen for efficiency. During the diffusion training stage, the same train-dev splits are used as in the adversarial training stage, thus ensuring no data leakage. At inference time, the number of ensembles is set to 10. Appendix C lists the hyper-parameters for each dataset.
5 Results and Analysis
5.1 Adversarial Robustness
| Dataset | PLM | Method | Clean% | AUA% | #Query | ||||
| TF | TB | BA | TF | TB | BA | ||||
| AGNews | BERT-base | Fine-Tuned | 94.4 | 10.2 | 25.4 | 27.1 | 348 | 372 | 379 |
| InfoBERT | 95.0 | 35.5 | 39.1 | 42.6 | 377 | 397 | 397 | ||
| FreeLB++ | 95.0 | 54.7 | 56.5 | 44.6 | 426 | 430 | 390 | ||
| EarlyRobust | 94.4 | 35.6 | 37.2 | 45.7 | 475 | 516 | 533 | ||
| RSMI | 94.3 | 52.6 | 56.7 | 55.4 | 680 | 737 | 687 | ||
| DiffuseDef-FreeLB++ (Ours) | 94.8 | 84.5 | 86.0 | 84.6 | 877 | 972 | 910 | ||
| DiffuseDef-RSMI (Ours) | 93.8 | 82.7 | 83.3 | 84.4 | 894 | 1029 | 930 | ||
| RoBERTa-base | Fine-Tuned | 94.9 | 34.1 | 36.9 | 43.6 | 372 | 396 | 410 | |
| InfoBERT | 95.5 | 40.2 | 45.2 | 48.6 | 392 | 421 | 430 | ||
| FreeLB++ | 95.4 | 57.5 | 62.9 | 55.9 | 444 | 467 | 447 | ||
| RSMI | 93.1 | 64.2 | 66.4 | 67.4 | 774 | 861 | 808 | ||
| DiffuseDef-FreeLB++ (Ours) | 95.3 | 85.6 | 87.6 | 85.3 | 880 | 976 | 906 | ||
| DiffuseDef-RSMI (Ours) | 92.9 | 82.9 | 83.5 | 82.2 | 905 | 925 | 1047 | ||
| IMDB | BERT-base | Fine-Tuned | 93.3 | 7.7 | 8.3 | 10.5 | 540 | 534 | 378 |
| InfoBERT | 93.9 | 29.2 | 25.4 | 30.7 | 642 | 644 | 390 | ||
| FreeLB++ | 94.3 | 44.2 | 39.6 | 40.6 | 784 | 829 | 426 | ||
| EarlyRobust | 92.7 | 49.7 | 46.8 | 43.8 | 2267 | 2788 | 1841 | ||
| RSMI | 90.9 | 60.0 | 54.4 | 51.1 | 2840 | 3455 | 2070 | ||
| DiffuseDef-FreeLB++ (Ours) | 94.4 | 82.1 | 83.0 | 84.0 | 3174 | 4348 | 2842 | ||
| DiffuseDef-RSMI (Ours) | 90.2 | 80.9 | 79.8 | 79.8 | 3590 | 4748 | 2901 | ||
| RoBERTa-base | Fine-Tuned | 94.6 | 21.3 | 17.9 | 13.6 | 587 | 671 | 493 | |
| InfoBERT | 94.8 | 30.9 | 27.9 | 21.8 | 681 | 760 | 549 | ||
| FreeLB++ | 95.3 | 46.0 | 42.1 | 33.9 | 829 | 974 | 637 | ||
| RSMI | 92.7 | 77.9 | 74.3 | 70.6 | 3443 | 4342 | 2619 | ||
| DiffuseDef-FreeLB++ (Ours) | 95.0 | 86.2 | 85.9 | 86.8 | 3573 | 4663 | 2941 | ||
| DiffuseDef-RSMI (Ours) | 92.4 | 84.7 | 84.1 | 84.3 | 3673 | 4782 | 3007 | ||
In Table 1, we compare the adversarial robustness of DiffuseDef with baselines and SOTA methods on AGNews and IMDB datasets trained with BERT and RoBERTa. DiffuseDef consistently outperforms all other methods on both datasets across both PLMs, exhibiting substantial improvements in accuracy under attack. After applying diffusion training, the AUA score for both FreeLB++ and RSMI models improves significantly, with an average increase of 30% AUA against the three attack methods. Note that despite the robust adversarial performance of the RSMI model, especially when trained with RoBERTa on the IMDB dataset, it still benefits from DiffuseDef. When comparing the clean accuracies to its base model (i.e. FreeLB++ and RSMI), DiffuseDef only shows a minor decline, between 0.2 and 0.7 accuracy score, which indicates that it can preserve the clean text performance while improving adversarial robustness. Moreover, models trained with DiffuseDef show a much smaller gap between clean accuracy and accuracy under attack, and such difference can be reduced to less than 10% AUA.
Another benefit of DiffuseDef is the increased number of adversarial queries needed to obtain a successful attack. Models applying DiffuseDef require over twice the number of queries on both datasets compared to the other methods. This increase is even larger on the IMDB dataset due to the longer text length. For example, DiffuseDef model requires on average over 3000 queries to achieve a successful attack while FreeLB++ only needs 400 to 800 queries. The substantial increase suggests that even if the attackers manage to construct a successful adversarial attack, they need 2x to 3x more time to find the attack on DiffuseDef than other models, affirming the improved robustness from diffusion training. In addition, we observe that the number of queries for denoising-based methods (i.e. RSMI, DiffuseDef) is generally higher than adversarial training-based methods (i.e. InfoBERT, FreeLB++). This is because denoising-based methods transform the hidden representations of the adversarial texts into a non-deterministic representation. The introduction of randomness in hidden states results in uncertainty in model logits, thus increasing the difficulty finding vulnerable words to attack (Wang et al., 2023).
5.2 Ablation - NLI Task
To understand how each component contributes to DiffuseDef, we conduct an ablation analysis on the QNLI dataset (Table 2). Compared to the fine-tuning baseline, FreeLB++ increases the AUA score from 21.5 to 45.6, showing the benefit of adversarial training. After applying diffusion training (with inference timestep ), the score is further improved to 49.2, showing that diffusion training complements adversarial training. Finally, ensembling enhances adversarial performance and improves the score to 66.7, with the number of queries growing from 392 to 485. Similar improvements in both AUA and number of queries is found with the RSMI model after applying diffusion training and ensembling, which validates that the two components are complementary and that DiffuseDef is compatible with multiple SOTA defense methods.
| Method | Clean% | AUA% | #Query |
| Fine-Tuned (BERT) | 90.8 | 21.5 | 195 |
| FreeLB++ | 90.3 | 45.6 | 253 |
| + diffusion training | 90.2 | 49.2 | 392 |
| + ensembling | 90.3 | 66.7 | 485 |
| RSMI | 87.4 | 35.2 | 314 |
| + diffusion training | 86.5 | 40.0 | 353 |
| + ensembling | 86.4 | 55.5 | 459 |
5.3 Robustness w.r.t Token Length
Figure 2 provides comparison of defense rate for different models by token length on the IMDB dataset. The defense rate is calculated as the percentage of test examples in which TextFooler fails to construct a successful attack. All models except RSMI show a consistent trend that the defense rate declines as the texts lengthen. This trend can be attributed to the nature of adversarial attacks as longer texts allow for the generation of more adversarial examples. Specifically, adversarial training defense methods like InfoBERT and FreeLB++ show poor performance on longer texts (more than 300 tokens), with the defense rate reduced to near 0. This drastic decline indicates that given an adequate number of queries, the attacker is guaranteed to find a successful attack to fool these models. Similarly, EarlyRobust exhibits a performance drop on long texts as it is based on FreeLB training. RSMI, however, performs worse on short texts, but its defense rate increases as the text length grows. Compared to all SOTA defense approaches, the two DiffuseDef variants show a more steadily declining trend and maintain a higher defense rate across all token lengths, i.e. DiffuseDef is more robust to input text length.
5.4 Effect of Additional Denoising Steps
In Figure 3 we study how the inference denoising steps can affect the adversarial performance. For the DiffuseDef model without ensembling, both AUA score and the number of queries required to attack increase as the inference denoising step is larger. As the denoising step grows from 1 to 30, the AUA score improves from 58 to 65 while the number of attack queries grows from 430 to 780. In contrast, for DiffuseDef with ensembling, the model maintains a stable but robust performance in AUA and number of queries, regardless of the increase of . Considering that the ensembling introduces a notable performance increase, the DiffuseDef model is likely to be hitting an upperbound in both metrics, thus no further improvement is reached by increasing the denoising steps. However, it also shows that with ensembling, DiffuseDef can be applied with a smaller for better efficiency while maintaining a robust adversarial performance.
5.5 Ensembling Diffused Hidden Representations
In DiffuseDef the text hidden state is diffused and ensembled to form a denoised hidden representation, which contributes significantly to the improved adversarial robustness. In this section, we study how the ensembling diffused hidden representation helps defend against adversarial attacks.
As mentioned in Section 2.1, attack methods need to first rank token importance based on its influence on prediction. Specifically, the importance score is calculated by comparing the change of model prediction probablities after removing each word. In Figure 4, we compare the distribution of max token importance score between FreeLB++ and its DiffuseDef counterpart. Both FreeLB++ and DiffuseDef show a long-tail distribution with over 80 percent examples having a max token importance score below 0.1. This suggests that in most cases changing one single token will not significantly alter the prediction for both models. However, DiffuseDef shows a notably lower percentage of tokens when the max importance score is between 0.9 and 1, where the attacker can easily find the vulnerable token to construct adversarial examples. This difference shows that DiffuseDef can complicate the process of important word searching, which accounts for the increased number of queries required for a successful attack.
| Method | L2 | Cosine |
| FreeLB++ | 12.53 | 0.35 |
| DiffuseDef-FreeLB++ | 10.66 | 0.27 |
| RSMI | 9.72 | 0.24 |
| DiffuseDef-RSMI | 8.61 | 0.21 |
In addition, DiffuseDef mitigates the difference between clean and adversarial texts by reducing the distance between their hidden states. In Table 3, we report the L2 and cosine distance between clean and adversarial hidden states for FreeLB++ and RSMI. Both show lower L2 and cosine distance after applying DiffuseDef, indicating that ensembling diffused representation repositions the adversarial example closer to the clean example, leading to the model maintaining its predictions.
5.6 Efficiency of DiffuseDef
| Method | Params | FLOPS |
| Fine-Tuned (BERT) | 110M | 46G |
| EarlyRobust | 82M | 32G |
| FreeLB++ | 110M | 46G |
| InfoBERT | 110M | 46G |
| RSMI | 110M | 92G |
| RanMask () | 110M | 459G |
| SAFER ( | 110M | 459G |
| DiffuseDef () | 120M | 96G |
| DiffuseDef () | 120M | 267G |
Given that DiffuseDef adds additional denoising and ensembling steps during inference, it inevitably increases the computation time compared to its base model. To study its efficiency, we report the number of model parameters and inference FLOPS in Table 4. In addition to the defense methods in Table 1, we also compare the efficiency of DiffuseDef with two other SOTA ensembling-based defense methods, i.e. RanMask (Zeng et al., 2023) and SAFER Ye et al. (2020).
All SOTA models have the same number of parameters as the fine-tuned BERT model, except EarlyRobust which applies attention head pruning for better efficiency. DiffuseDef, with 1 additional diffusion layer, increases the number of parameters from 110M to 120M. DiffuseDef requires more inference FLOPS than non ensembling-based baselines such as FreeLB++ and EarlyRobust. With and , the FLOPS for DiffuseDef doubles from 46G to 96G, nevertheless, this number is close to RSMI model (92G FLOPS) as it requires gradient information during inference. Despite this increase, DiffuseDef is more efficient than ensembling-based methods like RanMask and SAFER which need to go through a full forward pass for all ensembles. With the same ensembling number of 10, both RanMask and SAFER require 459G FLOPS, which is 10x the number for BERT baseline. In contrast, even with increased to 5, DiffuseDef can be run faster with 267G FLOPS, showing that it can mitigate the efficiency problem from ensembling while maintaining the benefit of improved robustness.
6 Conclusions
We propose a novel adversarial defense method, DiffuseDef, which combines adversarial training, diffusion training, and ensembling to improve model robustness to adversarial attacks. DiffuseDef can build on any existing adversarial training method, training an additional diffusion layer to predict and remove randomly sampled noise at a given timestep. During inference, the diffusion layer is used to denoise the adversarial hidden states, which are ensembled to construct a robust text representation. Our experiments validate the effectiveness and efficiency of DiffuseDef, which significantly outperforms SOTA on three common adversarial attack methods. Analysis shows that DiffuseDef makes it difficult to find vulnerable tokens to attack, and also reduces the difference between the hidden representations of clean and adversarial texts.
7 Limitations
Scope
Our experiments focus on defending against three common black-box adversarial attack methods, while whether DiffuseDef improves model robustness against white-box attacks is unclear. White-box attacks have access to model parameters and can utilize gradient information to construct adversarial examples more efficiently than black-box attacks. Defending against white-box attacks is more challenging, and we consider this as a future direction of DiffuseDef.
Comparison with additional approaches
Due to the length limit, we do not compare against all current approaches. However we do compare with the SOTA methods with best adversarial robustness based on our preliminary experiments.
Efficiency
Despite the fact that DiffuseDef is more efficient than existing ensembling-based methods, it still requires more model parameters and inference FLOPS than non-ensembling-based models to achieve a better robustness. Future directions of this work might involve efforts to reduce the size of diffusion layer and number of ensembles to make DiffuseDef more efficient.
8 Ethical Considerations
In this paper we propose a new method DiffuseDef which uses a diffusion layer as a denoiser to provide robust and efficient text representation. We demonstrate that the proposed method could significantly improve the robustness of NLP systems to adversarial attacks. However, DiffuseDef cannot defend against all adversarial attacks without limitations (e.g. number of perturbed words, semantic similarity between original and adversarial examples). Potential risks might include creation of new adversarial attacks devised specifically for DiffuseDef.
Appendix A Data Preparation
| Dataset | Train | Valid | Test | Avg Len |
| AGNews | 108K | 12K | 7K | 51.3 |
| IMDB | 40K | 5K | 5K | 311.9 |
| QNLI | 94K | 10K | 5K | 47.2 |
Table 5 presents the number of examples in train/valid/test splits and the average token length for the three datasets used in the experiments. For QNLI and AGNews datasets, we randomly split the training set into our train/valid splits, with a ratio of 0.9/0.1, and use its test set as our test split. For IMDB dataset, we randomly split the dataset into train/valid/test splits with a ratio of 0.8/0.1/0.1. All train/valid/test splitting is performed using a random seed of 42.
Appendix B Evaluation Constraints
| Dataset | |||
| AGNews | 0.84 | 50 | 0.3 |
| IMDB | 0.84 | 50 | 0.1 |
| QNLI | 0.84 | 50 | 0.2 |
When evaluating with adversarial attack, We follow the parameter settings for TextAttack as suggested in (Li et al., 2021). The minimum semantic similarity between the clean text and adversarial text is set to 0.84, with the score computed using Universal Sentence Encoder (Cer et al., 2018). The maximum number of candidate substitution from attacker is 50, thus the maximum number of queries where is the number of tokens. Finally, the maximum percentage of changed tokens is set to 0.3/0.1/0.2 for AGNews, IMDB, and QNLI dataset respectively.
Appendix C Training
| AGNews | IMDB | QNLI | |
| Epochs | 100 | 100 | 100 |
| Batch size | 64 | 64 | 64 |
| Sequence len | 128 | 256 | 256 |
| Dropout | 0.1 | 0.1 | 0.1 |
| Optimizer | AdamW | AdamW | AdamW |
| Lr | 2e-5 | 2e-5 | 2e-5 |
| 30 | 10 | 30 | |
| 5 | 5 | 5 | |
| 10 | 10 | 10 |
The details on hyper-parameters of diffusion training can be found in Table 7. All models are trained on a single RTX A6000 GPU. The diffusion training of 100 epochs takes 6/4/3 hours on AGNews, IMDB, QNLI datasets respectively.
Appendix D License for Scientific Artifacts
| Artifact | License |
| AGNews (Zhang et al., 2015b) | Custom (non-commercial) |
| IMDB (Maas et al., 2011b) | - |
| QNLI (Wang et al., 2018) | CC BY-SA 4.0 |
| transformers (Wolf et al., 2020) | Apache License 2.0 |
| TextAttack (Morris et al., 2020) | MIT License |
| BERT (Devlin et al., 2019) | Apache License 2.0 |
| RoBERTa (Liu et al., 2019) | MIT License |
Table 8 lists the scientific artifacts including data, codes, and models used in this paper. The use of these artifacts in this paper is consistent with their intended use, i.e. for scientific research only. The data used in the experiment is in English and does not contain personally identifying info or offensive content.
Appendix E Example of noising and denoising in DiffuseDef
Adding and removing noise to hidden states are essential features in DiffuseDef which contribute to the improved adversarial robustness. To study how adding or removing noise can affect the semantic meaning of the text, we feed the hidden states to the pretrained BERT model with masked language modeling (MLM) head to generate the text output.
In Table 9, we present the MLM outputs from hidden states added with different steps of noise and the MLM outputs from noise hidden states denoised with same number of steps. In the example shown, with more noise added some semantic information can be lost and replaced by symbols or function words like "." or "the". In contrast, denoising for the same number of steps help alleviate such information lost. For example, the word "IBM" can be recovered from the noise.
However, in practise it is not possible to assume number of denoising steps therefore in Table 10 we show the MLM outputs of denoised hidden states directly from clean and adversarial texts. On clean text, we observe that a higher number of denoising steps can result in more abstraction of the texts. For example, more words are replaced with "the" in the MLM outputs as grows. However, words related to the topic (e.g. "Manchester United", "Liverpool") are kept during the denoising process, thus the model can predict correctly. Similarly, the trend of abstraction can be also found on adversarial text while we observe that the denoising can help remove the adversarial noise / perturbation and recover the word "united" from "nation", thus resulting its correct prediction on the adversarial text.
| MLM Output (add noise) | MLM Output (add noise then denoise) | |
| 0 | IBM Chips May Someday Heal Themselves New technology applies electrical fuses to help identify and repair faults. | - |
| 5 | the ibm chips may someday heal themselves new technology introduces electrical fuses to help identify and repair faults. | the ibm chips may someday heal themselves new technology introduces electrical fuses to help identify and repair faults. |
| 6 | ) ibm chips may someday heal themselves new technology introduces electrical fuses to help identify and repair faults. | the ibm chips may someday heal themselves new technology introduces electrical fuses to help identify and repair faults. |
| 7 | the. chips may someday heal themselves new technology introduces electrical fuses to help identify and repair faults. | the. chips may someday heal themselves new technology uses electrical fuses to help identify and repair faults. |
| 8 | ).. may someday heal themselves new technology introduces electrical fuses to help identify and repair faults. | the ibm. may someday heal themselves new technology uses electrical fuses to help identify and repair faults. |
| 9 | the ibm chips may someday heal themselves new technology uses electrical fuses to help identify and repair faults. | the ibm chips may someday heal themselves new technology introduces electrical fuses to help identify and repair faults. |
| 10 | the. chips may someday heal themselves new technology extends electrical fuses to help identify and repair faults. | the ibm. may someday heal themselves new technology develops electrical fuses to help identify and repair faults. |
| Clean Text / MLM Output | Adv Text / MLM Output | Pred clean | Pred adv | |
| 0 | United Apology over Website Abuse Manchester United have been forced to issue an embarrassing apology to Liverpool for an ill-advised attack on the Anfield outfit on its own website. | United Apology over Website Abuse Manchester Nations have been forced to issue an embarrassing apology to Liverpool for an ill-advised attack on the Anfield outfit on its own website. | Sports | World |
| 1 | football. apology over website abuse manchester united have been - to issue an embarrassing apology to liverpool for an the - advised attack on the anfield outfit on its own website. | the. apology over website abuse manchester nations have been the to issue an embarrassing apology to liverpool for an the - advised attack on the anfield outfit on its own website. | Sports | World |
| 2 | the. apology over website abuse manchester united have been - to issue an embarrassing apology to liverpool for an the - advised attack on the anfield outfit on its own website. | the. apology over website abuse manchester nations have been the to issue an embarrassing apology to liverpool for an the - advised attack on the anfield outfit on its own website. | Sports | World |
| 3 | the. apology over website abuse manchester united have been the to issue an embarrassing apology to liverpool for an the - advised attack on the anfield outfit on its own website. | the. apology over website abuse manchester s have been the to issue an embarrassing apology to liverpool for an the - advised attack on the anfield outfit on its own website. | Sports | Sports |
| 4 | the. apology over website abuse manchester united have the - to issue an embarrassing apology to liverpool for an the - advised attack on the anfield outfit on its own website. | the. apology over website abuse manchester s have been the to issue an embarrassing apology to liverpool for an the - advised attack on the anfield outfit on its own website. | Sports | Sports |
| 5 | the. apology over website abuse manchester united have the the to issue an a apology to liverpool for an the - advised attack on the anfield outfit on its own website. | the. apology over website abuse manchester united have been the to issue an the apology to liverpool for an’- advised attack on the anfield outfit on its own website. | Sports | Sports |
Appendix F Confusion Matrix under Attack
Figure 5 and 6 present the confusion matrixes of models prediction on clean text and on adversarial texts (successful attack example) on AGNews and IMDB test sets respectively.

### v2

@@
-DiffuseDef: Improved Robustness to Adversarial Attacks
+DiffuseDef: Improved Robustness to Adversarial Attacks via Iterative Denoising
 Abstract
 Pretrained language models have significantly advanced performance across various natural language processing tasks.
-However, adversarial attacks continue to pose a critical challenge to system built using these models, as they can be exploited with carefully crafted adversarial texts.
-Inspired by the ability of diffusion models to predict and reduce noise in computer vision, we propose a novel and flexible adversarial defense method for language classification tasks, DiffuseDef111https://github.com/Nickeilf/DiffuseDef, which incorporates a diffusion layer as a denoiser between the encoder and the classifier.
+However, adversarial attacks continue to pose a critical challenge to systems built using these models, as they can be exploited with carefully crafted adversarial texts.
+Inspired by the ability of diffusion models to predict and reduce noise in computer vision, we propose a novel and flexible adversarial defense method for language classification tasks, DiffuseDef 111Codes available at: https://github.com/Nickeilf/DiffuseDef, which incorporates a diffusion layer as a denoiser between the encoder and the classifier.
+The diffusion layer is trained on top of the existing classifier, ensuring seamless integration with any model in a plug-and-play manner.
 During inference, the adversarial hidden state is first combined with sampled noise, then denoised iteratively and finally ensembled to produce a robust text representation.
-By integrating adversarial training, denoising, and ensembling techniques, we show that DiffuseDef improves over different existing adversarial defense methods and achieves state-of-the-art performance against common adversarial attacks.
-DiffuseDef: Improved Robustness to Adversarial Attacks
-Zhenhao Li, Marek Rei, Lucia Specia Language and Multimodal AI (LAMA) Lab, Imperial College London {zhenhao.li18, marek.rei, l.specia}@imperial.ac.uk
+By integrating adversarial training, denoising, and ensembling techniques, we show that DiffuseDef improves over existing adversarial defense methods and achieves state-of-the-art performance against common black-box and white-box adversarial attacks.
+DiffuseDef: Improved Robustness to Adversarial Attacks via Iterative Denoising
+Zhenhao Li, Huichi Zhou, Marek Rei, Lucia Specia Language and Multimodal AI (LAMA) Lab, Imperial College London {zhenhao.li18, h.zhou24, marek.rei, l.specia}@imperial.ac.uk
 1 Introduction
@@
 In order to improve robustness to adversarial attacks, various defense methods have been proposed, such as adversarial training (Zhu et al., 2020; Si et al., 2021; Zhou et al., 2021; Xi et al., 2022), text denoising (Nguyen Minh and Luu, 2022; Wang et al., 2023), ensembling (Zhou et al., 2021; Zeng et al., 2023; Li et al., 2023), etc.
-However, existing defense methods either assume the test-time perturbation/attack set is similar to that used in training (Li et al., 2021), or are limited to specific architectures (Xi et al., 2022), or at inference time require large computational cost, thereby limiting their practical applicability.
+However, existing defense methods either assume the test-time perturbation/attack set is similar to that used in training (Li et al., 2021), or are limited to specific architectures (Xi et al., 2022), or at inference time require large computational costs, thereby limiting their practical applicability.
 Diffusion models are commonly used in computer vision (CV) to generate high-quality images by predicting and removing noise from a sampled noisy image.
-Therefore, they can be adopted to remove noise from adversarial images and thus improve robustness to attacks (Nie et al., 2022).
+Therefore, they can be adopted to remove noise from adversarial images and thus improve robustness to attacks (Im et al., 2016; Nie et al., 2022; Gulsen et al., 2024).
 However, in NLP very limited research has investigated adversarial defense with diffusion models due to the discrete and contextual nature of text data.
@@
 (2023) and Yuan et al.
-(2024) which apply diffusion on texts or labels, DiffuseDef directly removes noise to the hidden states, providing a more effective and robust text representation to defend against adversarial texts.
-Compared to diffusion-based defense in CV (Nie et al., 2022), DiffuseDef further enhances robustness with ensembling and improves efficiency with fewer diffusion steps.
+(2024), which apply diffusion on texts or labels, DiffuseDef directly removes noise from the hidden states, providing a more effective and robust text representation to defend against adversarial texts.
+Different from diffusion-based defenses in CV (Nie et al., 2022; Gulsen et al., 2024) that apply purification to input images, DiffuseDef only performs denoising at the final hidden layer, resulting in an improvement in efficiency.
+This improvement is further enhanced when combined with ensembling, as it eliminates the need for a forward pass through all classifier parameters for each ensemble.
 DiffuseDef combines adversarial training with diffusion training, where the diffusion layer is trained to predict randomly sampled noise at a given timestep.
 During inference, the diffusion layer serves as a denoiser, iteratively removing noise from adversarial hidden states to yield a robust hidden representation.
-Moreover, we adopt the ensembling strategy by first adding random noise to text hidden states to create multiple variants then denoising them via the diffusion layer.
+Moreover, we adopt the ensembling strategy by first adding random noise to the text hidden states to create multiple variants, then denoising them via the diffusion layer.
 The model output is made by averaging all denoised hidden states.
 Since ensembling happens solely at the diffusion layer, DiffuseDef is more efficient than traditional ensembling-based methods (Ye et al., 2020; Zeng et al., 2023), which require a full forward pass through all model parameters.
-Through systematic experimentation, we demonstrate that DiffuseDef outperforms strong defense methods and is able to defend against multiple types of adversarial attacks, while preserving performance on clean texts.
-Our analysis also reveals that the ensembling diffused representations provides a stronger defense against finding vulnerable words to attack and can reduce the distance in latent space between adversarial texts and their clean text counterpart.
-Our contributions can be summarised as follows:
+Through systematic experimentation, we demonstrate that DiffuseDef outperforms strong defense methods and is able to defend against multiple types of black-box and white-box adversarial attacks, while preserving performance on clean texts.
+Our analysis also reveals that the ensembling diffused representation provides a stronger defense against finding vulnerable words to attack and can reduce the distance in latent space between adversarial texts and their clean text counterparts.
+Our contributions can be summarized as follows:
 We propose DiffuseDef, a novel and flexible adversarial defense method that can be added on top of any existing adversarial defense methods to further improve robustness to adversarial attacks.
-DiffuseDef outperforms existing adversarial methods and achieves state-of-the-art performance against prevalent adversarial attacks.
+DiffuseDef outperforms existing adversarial methods and achieves state-of-the-art performance against prevalent black-box and white-box adversarial attacks.
 Through extensive analysis, we demonstrate the effectiveness of the ensembling diffused representation and the efficiency of DiffuseDef compared to existing ensembling-based methods.
@@
 2.1 Textual Adversarial Attacks
-Textual adversarial attacks focus on constructing adversarial examples from an original text that maximise the likelihood of incorrect predictions by a neural network.
-These attacks require adversarial examples to be perceptually similar to the original text, which is typically achieved by introducing subtle perturbations to the original text, such as character swapping (Gao et al., 2018; Ebrahimi et al., 2018), synonym-substitutions (Ren et al., 2019; Yoo and Qi, 2021), and paraphrasing (Gan and Ng, 2019; Huang and Chang, 2021).
+Textual adversarial attacks focus on constructing adversarial examples from an original text that maximize the likelihood of incorrect predictions by a neural network.
+These attacks require adversarial examples to be perceptually similar to the original text, which is typically achieved by introducing subtle perturbations to the original text, such as character swapping (Gao et al., 2018; Ebrahimi et al., 2018), synonym-substitutions (Ren et al., 2019; Yoo and Qi, 2021; Liu et al., 2023), and paraphrasing (Gan and Ng, 2019; Huang and Chang, 2021).
 Taking the text classification task as an example, given a classifier that maps an input sequence of words to its designated label , the goal of the attack model is to construct an adversarial example to fool the classifier, where is a subtle adversarial perturbation constrained by .
@@
 The substitution process is guided by various constraints to ensure the adversarial example remains natural and semantically equivalent to the original text.
-Common constraints include thresholding the similarity between the replacement word embedding and the substituted word embedding, or ensuring the semantic similarity between sentence vectors modeled from Universal Sentence Encoder (Cer et al., 2018).
+Common constraints include thresholding the similarity between the replacement word embedding and the substituted word embedding, or ensuring the semantic similarity (Zhou et al., 2024) between sentence vectors modeled from Universal Sentence Encoder (Cer et al., 2018).
 Despite these constraints, current textual adversarial attacks still pose significant challenges to NLP models (Liu et al., 2022; Xu et al., 2021; Yuan et al., 2023), highlighting the necessity for defense methods for better adversarial robustness.
@@
 They can be grouped into three categories: training-based, ensembling-based, and denoising-based methods.
-Adversarial training improves the robustness of the model to adversarial examples through strategies like data augmentation (Si et al., 2021) and adversarial regularisation (Madry et al., 2018; Zhu et al., 2020; Wang et al., 2021; Xi et al., 2022; Gao et al., 2023).
-However, adversarial training methods are limited as they assumes similar train-test adversarial examples, and thus tend to overfit to specific types of adversarial attacks.
+Adversarial training improves the robustness of the model to adversarial examples through strategies like data augmentation (Si et al., 2021; Bao et al., 2021) and adversarial regularization (Madry et al., 2018; Zhu et al., 2020; Wang et al., 2021; Xi et al., 2022; Gao et al., 2023; Formento et al., 2024).
+However, adversarial training methods are limited as they assume similar train-test adversarial examples, and thus tend to overfit to specific types of adversarial attacks.
 Ensembling-based methods generate multiple variants of the input text at inference time and ensemble model predictions over all the variants (Ye et al., 2020; Zhou et al., 2021; Zeng et al., 2023; Li et al., 2023), but they can be inefficient given that model predictions are needed on every ensemble, increasing the inference time with the number of ensembles.
-More recently, denoising-based methods have been proposed to improve adversarial robustness by mapping the vector representation of the adversarial text to another point in the latent space that is close to the clean text (Nguyen Minh and Luu, 2022; Wang et al., 2023; Moon et al., 2023; Yuan et al., 2024).
+More recently, denoising-based methods have been proposed to improve adversarial robustness by mapping the vector representation of the adversarial text to another point in the latent space that is close to the clean text (Nguyen Minh and Luu, 2022; Wang et al., 2023; Moon et al., 2023; Yuan et al., 2024; Ji et al., 2024).
 The denoised representation makes it more difficult to find vulnerable words to attack, thus improving adversarial robustness (Wang et al., 2023).
@@
 The training process is split into two stages: adversarial training and diffusion training (Figure 1).
-The adversarial training stage employs any neural network-based adversarial training methods like FreeLB++ (Li et al., 2021) and RSMI (Moon et al., 2023), which optimise the encoder and classifier for robustness by perturbing the latent representation of the text input.
+The adversarial training stage employs a neural network-based adversarial training method like FreeLB++ (Li et al., 2021) or RSMI (Moon et al., 2023), which optimizes the encoder and classifier for robustness by perturbing the latent representation of the text input.
 In the diffusion training stage, only the diffusion layer is trained to predict random noise added to the clean text hidden state at different timesteps, enabling it to denoise the adversarial hidden state at inference time.
-The pretrained encoder, however is frozen during this stage.
+The pretrained encoder, however, is frozen during this stage.
 Since the pretrained encoder is only used for feature extraction, the diffusion training method is compatible with any neural network-based adversarial training method.
 Given an input sequence of tokens , the pretrained encoder extracts the hidden state .
-A random Gaussian noise is sampled to perturb hidden state .
+A random Gaussian noise is sampled to perturb the hidden state .
 Sohl-Dickstein et al.
@@
 The time embedding receives the diffusion timestep as input and produces an embedding , which is added to as input for the transformer layer.
-Finally, the transformer layer outputs the predicted noise , and mean square error is used to compute the loss between predicted noise and actual sampled noise .
+Finally, the transformer layer outputs the predicted noise , and mean square error (MSE) is used to compute the loss between the predicted noise and the actual sampled noise .
 3.2 Inference
-Leveraging the diffusion layer’s ability to predict noise at a given timestep , we utilise it as a denoiser during inference by iteratively performing the reverse diffusion steps, which sample from to produce the denoised hidden state
+Leveraging the diffusion layer’s ability to predict noise at a given timestep , we utilize it as a denoiser during inference by iteratively performing the reverse diffusion steps, which sample from to produce the denoised hidden state
 where is the predicted noise from diffusion layer and . The denoised hidden state can thus be computed with
@@
 Unlike Ho et al.
-(2020) where the reverse diffusion step starts with pure noise sampled from standard normal distribution, we assume the noisy hidden state is already an intermediate state in the reverse diffusion steps.
+(2020), where the reverse diffusion step starts with pure noise sampled from a standard normal distribution, we assume the noisy hidden state is already an intermediate state in the reverse diffusion steps.
 This allows us to use a smaller number of than the training timestep to prevent the denoised hidden states from diverging substantially from the initial hidden state .
 This sequence of denoising steps creates the final denoised hidden states , which are averaged and used by the classifier to output the final predicted label.
-This process is summarised in Algorithm 1.
+This process is summarized in Algorithm 1.
 4 Experiments
-Datasets
+Datasets.
 We focus on two common NLP tasks in our experiments: topic classification and natural language inference (NLI).
-In the text classification task, we compare our method with other defense algorithms on two standard datasets for adversarial defense: AG News (Zhang et al., 2015a) and IMDB (Maas et al., 2011a) datasets.
-In the NLI task, we perform an ablation analysis with the Question-answering NLI (QNLI) dataset (Wang et al., 2018).
+We conduct experiments on three common datasets: AG News (Zhang et al., 2015a), IMDB (Maas et al., 2011a), and QNLI datasets (Wang et al., 2018).
 We randomly split AGNews, IMDB, and QNLI datasets into train, validation, and test splits.
-Evaluation
-Following previous work on adversarial defense, we use three benchmarking attack methods to evaluate the robustness of DiffuseDef: TextFooler (TF) (Jin et al., 2020), TextBugger (TB) (Li et al., 2019), and Bert-Attack (BA) (Li et al., 2020).
-The three attack methods create adversarial attacks in different granularities: character-level perturbation (TextBugger), word substitution (TextFooler), and subword substitution (BertAttack).
+Appendix A gives details on data preparation.
+Evaluation.
+Following previous work on adversarial defense, we use three benchmarking black-box attack methods to evaluate the robustness of DiffuseDef: TextFooler (TF) (Jin et al., 2020), TextBugger (TB) (Li et al., 2019), and Bert-Attack (BA) (Li et al., 2020).
+We also evaluated the robustness against two white-box attacks: T-PGD (Yuan et al., 2023) and SemAttack (Wang et al., 2022).
 Regarding evaluation metrics, we measure the clean accuracy (Clean%) on the test set, the accuracy under attack (AUA%), and the number of adversarial queries (#Query) needed for a successful attack.
@@
 4.1 Comparison to SOTA
-We compare our proposed method with state-of-the-art adversarial defense approaches, trained using both BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) as backbones: Fine-tune: Fine-tuning pretrained models on downstream task with no defense method applied222”Fine-tune” is a baseline approach used to illustrate the effect of adversarial attacks..
-InfoBERT (Wang et al., 2021): Applying mutual-information-based regularizers during fine-tuning of pretrained models to improve robustness.
+We compare our proposed method with state-of-the-art adversarial defense approaches, trained using both BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) as backbones: Fine-tune: Fine-tuning pretrained models on downstream tasks with no defense method applied222“Fine-tuned” is a baseline approach used to illustrate the effect of adversarial attacks..
+InfoBERT (Wang et al., 2021): Applying mutual-information-based regularizer during fine-tuning of pretrained models to improve robustness.
 FreeLB++ (Li et al., 2021): An adversarial training method improving on FreeLB(Zhu et al., 2020), which adds adversarial perturbations to word embedding during fine-tuning.
@@
 (Xi et al., 2022): Extracting early-bird subnetworks and pruning pretrained models for efficient adversarial training.
-RSMI (Moon et al., 2023): A two-stage training method that combines randomised smoothing and masked inference to improve adversarial robustness.
+RMLM (Wang et al., 2023): A denoising-based model combined with adversarial text detection.
+ATINTER (Gupta et al., 2023): A fine-tuned T5 model that rewrites the input adversarial texts, and the classifier predicts labels from the rewritten texts.
+RSMI (Moon et al., 2023): A two-stage training method that combines randomized smoothing and masked inference to improve adversarial robustness.
 4.2 Implementation and Settings
@@
 In the diffusion layer, only one transformer encoder layer (Vaswani et al., 2017) is used.
-The maximum noising timestep during training is set to 30 for AGNews and QNLI datasets, and 10 for IMDB dataset, while at inference time, we only apply 5 denoising steps for .
+The maximum noising timestep during training is set to 30 for AGNews and QNLI datasets, and 10 for the IMDB dataset, while at inference time, we only apply 5 denoising steps for .
 We follow (Ho et al., 2020) to use a linear schedule from to .
-The diffusion layer is trained for 100 epochs, with the base classifier parameters frozen for efficiency.
+The base encoder and classifier are first fine-tuned for 10 epochs, and the diffusion layer is trained for 100 epochs, with the base encoder and classifier parameters frozen for efficiency.
 During the diffusion training stage, the same train-dev splits are used as in the adversarial training stage, thus ensuring no data leakage.
 At inference time, the number of ensembles is set to 10.
-Appendix C lists the hyper-parameters for each dataset.
+Appendix C lists the training hyperparameters.
 5 Results and Analysis
 5.1 Adversarial Robustness
-| Dataset | PLM | Method | Clean% | AUA% | #Query | ||||
-| TF | TB | BA | TF | TB | BA | ||||
-| AGNews | BERT-base | Fine-Tuned | 94.4 | 10.2 | 25.4 | 27.1 | 348 | 372 | 379 |
-| InfoBERT | 95.0 | 35.5 | 39.1 | 42.6 | 377 | 397 | 397 | ||
-| FreeLB++ | 95.0 | 54.7 | 56.5 | 44.6 | 426 | 430 | 390 | ||
-| EarlyRobust | 94.4 | 35.6 | 37.2 | 45.7 | 475 | 516 | 533 | ||
-| RSMI | 94.3 | 52.6 | 56.7 | 55.4 | 680 | 737 | 687 | ||
-| DiffuseDef-FreeLB++ (Ours) | 94.8 | 84.5 | 86.0 | 84.6 | 877 | 972 | 910 | ||
-| DiffuseDef-RSMI (Ours) | 93.8 | 82.7 | 83.3 | 84.4 | 894 | 1029 | 930 | ||
-| RoBERTa-base | Fine-Tuned | 94.9 | 34.1 | 36.9 | 43.6 | 372 | 396 | 410 | |
-| InfoBERT | 95.5 | 40.2 | 45.2 | 48.6 | 392 | 421 | 430 | ||
-| FreeLB++ | 95.4 | 57.5 | 62.9 | 55.9 | 444 | 467 | 447 | ||
-| RSMI | 93.1 | 64.2 | 66.4 | 67.4 | 774 | 861 | 808 | ||
-| DiffuseDef-FreeLB++ (Ours) | 95.3 | 85.6 | 87.6 | 85.3 | 880 | 976 | 906 | ||
-| DiffuseDef-RSMI (Ours) | 92.9 | 82.9 | 83.5 | 82.2 | 905 | 925 | 1047 | ||
-| IMDB | BERT-base | Fine-Tuned | 93.3 | 7.7 | 8.3 | 10.5 | 540 | 534 | 378 |
-| InfoBERT | 93.9 | 29.2 | 25.4 | 30.7 | 642 | 644 | 390 | ||
-| FreeLB++ | 94.3 | 44.2 | 39.6 | 40.6 | 784 | 829 | 426 | ||
-| EarlyRobust | 92.7 | 49.7 | 46.8 | 43.8 | 2267 | 2788 | 1841 | ||
-| RSMI | 90.9 | 60.0 | 54.4 | 51.1 | 2840 | 3455 | 2070 | ||
-| DiffuseDef-FreeLB++ (Ours) | 94.4 | 82.1 | 83.0 | 84.0 | 3174 | 4348 | 2842 | ||
-| DiffuseDef-RSMI (Ours) | 90.2 | 80.9 | 79.8 | 79.8 | 3590 | 4748 | 2901 | ||
-| RoBERTa-base | Fine-Tuned | 94.6 | 21.3 | 17.9 | 13.6 | 587 | 671 | 493 | |
-| InfoBERT | 94.8 | 30.9 | 27.9 | 21.8 | 681 | 760 | 549 | ||
-| FreeLB++ | 95.3 | 46.0 | 42.1 | 33.9 | 829 | 974 | 637 | ||
-| RSMI | 92.7 | 77.9 | 74.3 | 70.6 | 3443 | 4342 | 2619 | ||
-| DiffuseDef-FreeLB++ (Ours) | 95.0 | 86.2 | 85.9 | 86.8 | 3573 | 4663 | 2941 | ||
-| DiffuseDef-RSMI (Ours) | 92.4 | 84.7 | 84.1 | 84.3 | 3673 | 4782 | 3007 | ||
-In Table 1, we compare the adversarial robustness of DiffuseDef with baselines and SOTA methods on AGNews and IMDB datasets trained with BERT and RoBERTa.
-DiffuseDef consistently outperforms all other methods on both datasets across both PLMs, exhibiting substantial improvements in accuracy under attack.
+| Dataset | Method | Clean% | AUA% | #Query | ||||
+| TF | TB | BA | TF | TB | BA | |||
+| AGNews | Fine-Tuned | 94.4 | 10.2 | 25.4 | 27.1 | 348 | 372 | 379 |
+| InfoBERT (Wang et al., 2021) | 95.0 | 35.5 | 39.1 | 42.6 | 377 | 397 | 397 | |
+| FreeLB++ (Li et al., 2021) | 95.0 | 54.7 | 56.5 | 44.6 | 426 | 430 | 390 | |
+| EarlyRobust (Xi et al., 2022) | 94.4 | 35.6 | 37.2 | 45.7 | 475 | 516 | 533 | |
+| RMLM (Wang et al., 2023) | 94.3 | 45.6 | 44.4 | 54.1 | 642 | 738 | 705 | |
+| ATINTER (Gupta et al., 2023) | 94.2 | 68.0 | 59.0 | 81.0 | 527 | 235 | 122 | |
+| RSMI (Moon et al., 2023) | 94.3 | 52.6 | 56.7 | 55.4 | 680 | 737 | 687 | |
+| DiffuseDef-FreeLB++ (Ours) | 94.8 | 84.5 | 86.0 | 84.6 | 877 | 972 | 910 | |
+| DiffuseDef-RSMI (Ours) | 93.8 | 82.7 | 83.3 | 84.4 | 894 | 1029 | 930 | |
+| IMDB | Fine-Tuned | 93.3 | 7.7 | 8.3 | 10.5 | 540 | 534 | 378 |
+| InfoBERT (Wang et al., 2021) | 93.9 | 29.2 | 25.4 | 30.7 | 642 | 644 | 390 | |
+| FreeLB++ (Li et al., 2021) | 94.3 | 44.2 | 39.6 | 40.6 | 784 | 829 | 426 | |
+| EarlyRobust (Xi et al., 2022) | 92.7 | 49.7 | 46.8 | 43.8 | 2267 | 2788 | 1841 | |
+| RMLM (Wang et al., 2023) | 93.6 | 48.8 | 47.9 | 47.1 | 2554 | 3258 | 2064 | |
+| ATINTER (Gupta et al., 2023) | 94.3 | 25.0 | 15.0 | 39.0 | 1173 | 645 | 637 | |
+| RSMI (Moon et al., 2023) | 90.9 | 60.0 | 54.4 | 51.1 | 2840 | 3455 | 2070 | |
+| DiffuseDef-FreeLB++ (Ours) | 94.4 | 82.1 | 83.0 | 84.0 | 3174 | 4348 | 2842 | |
+| DiffuseDef-RSMI (Ours) | 90.2 | 80.9 | 79.8 | 79.8 | 3590 | 4748 | 2901 | |
+| QNLI | Fine-Tuned | 90.8 | 21.5 | 15.5 | 13.3 | 195 | 206 | 177 |
+| InfoBERT (Wang et al., 2021) | 91.2 | 27.8 | 22.8 | 18.7 | 217 | 232 | 201 | |
+| FreeLB++ (Li et al., 2021) | 90.3 | 45.6 | 40.2 | 30.5 | 253 | 279 | 226 | |
+| EarlyRobust (Xi et al., 2022) | 89.2 | 24.3 | 21.2 | 19.1 | 265 | 292 | 255 | |
+| ATINTER (Gupta et al., 2023) | 90.4 | 43.0 | 24.8 | 34.0 | 185 | 115 | 89 | |
+| RSMI (Moon et al., 2023) | 87.4 | 35.2 | 30.9 | 28.2 | 314 | 353 | 314 | |
+| DiffuseDef-FreeLB++ (Ours) | 90.3 | 66.7 | 65.3 | 64.4 | 485 | 587 | 543 | |
+| DiffuseDef-RSMI (Ours) | 86.4 | 55.5 | 54.8 | 57.2 | 459 | 569 | 528 | |
+In Table 1, we compare the adversarial robustness of DiffuseDef with baselines and SOTA methods on AGNews, IMDB, and QNLI datasets trained with BERT (results for RoBERTa in Appendix E).
+DiffuseDef consistently outperforms all other methods on all three datasets across the three attacks, exhibiting substantial improvements in accuracy under attack.
 After applying diffusion training, the AUA score for both FreeLB++ and RSMI models improves significantly, with an average increase of 30% AUA against the three attack methods.
-Note that despite the robust adversarial performance of the RSMI model, especially when trained with RoBERTa on the IMDB dataset, it still benefits from DiffuseDef.
-When comparing the clean accuracies to its base model (i.e.
-FreeLB++ and RSMI), DiffuseDef only shows a minor decline, between 0.2 and 0.7 accuracy score, which indicates that it can preserve the clean text performance while improving adversarial robustness.
-Moreover, models trained with DiffuseDef show a much smaller gap between clean accuracy and accuracy under attack, and such difference can be reduced to less than 10% AUA.
+When comparing the clean accuracies to its base model (i.e., FreeLB++ and RSMI), DiffuseDef only shows a minor decline, between 0.2 and 0.7 accuracy score, which indicates that it can preserve the clean text performance while improving adversarial robustness.
+Moreover, models trained with DiffuseDef show a much smaller gap between clean accuracy and accuracy under attack, and such a difference can be reduced to less than 10% AUA.
 Another benefit of DiffuseDef is the increased number of adversarial queries needed to obtain a successful attack.
-Models applying DiffuseDef require over twice the number of queries on both datasets compared to the other methods.
+Models applying DiffuseDef require over twice the number of queries on all three datasets compared to the other methods.
 This increase is even larger on the IMDB dataset due to the longer text length.
-For example, DiffuseDef model requires on average over 3000 queries to achieve a successful attack while FreeLB++ only needs 400 to 800 queries.
+For example, the DiffuseDef model requires on average over 3000 queries to achieve a successful attack, while FreeLB++ can only defend for 400 to 800 queries.
 The substantial increase suggests that even if the attackers manage to construct a successful adversarial attack, they need 2x to 3x more time to find the attack on DiffuseDef than other models, affirming the improved robustness from diffusion training.
-In addition, we observe that the number of queries for denoising-based methods (i.e.
-RSMI, DiffuseDef) is generally higher than adversarial training-based methods (i.e.
-InfoBERT, FreeLB++).
+In addition, we observe that the number of queries for denoising-based methods (i.e., RSMI, RMLM, and DiffuseDef) is generally higher than adversarial training-based methods (i.e., InfoBERT, FreeLB++).
 This is because denoising-based methods transform the hidden representations of the adversarial texts into a non-deterministic representation.
-The introduction of randomness in hidden states results in uncertainty in model logits, thus increasing the difficulty finding vulnerable words to attack (Wang et al., 2023).
-5.2 Ablation - NLI Task
-To understand how each component contributes to DiffuseDef, we conduct an ablation analysis on the QNLI dataset (Table 2).
-Compared to the fine-tuning baseline, FreeLB++ increases the AUA score from 21.5 to 45.6, showing the benefit of adversarial training.
-After applying diffusion training (with inference timestep ), the score is further improved to 49.2, showing that diffusion training complements adversarial training.
-Finally, ensembling enhances adversarial performance and improves the score to 66.7, with the number of queries growing from 392 to 485.
-Similar improvements in both AUA and number of queries is found with the RSMI model after applying diffusion training and ensembling, which validates that the two components are complementary and that DiffuseDef is compatible with multiple SOTA defense methods.
-| Method | Clean% | AUA% | #Query |
-| Fine-Tuned (BERT) | 90.8 | 21.5 | 195 |
-| FreeLB++ | 90.3 | 45.6 | 253 |
-| + diffusion training | 90.2 | 49.2 | 392 |
-| + ensembling | 90.3 | 66.7 | 485 |
-| RSMI | 87.4 | 35.2 | 314 |
-| + diffusion training | 86.5 | 40.0 | 353 |
-| + ensembling | 86.4 | 55.5 | 459 |
-5.3 Robustness w.r.t Token Length
-Figure 2 provides comparison of defense rate for different models by token length on the IMDB dataset.
-The defense rate is calculated as the percentage of test examples in which TextFooler fails to construct a successful attack.
-All models except RSMI show a consistent trend that the defense rate declines as the texts lengthen.
-This trend can be attributed to the nature of adversarial attacks as longer texts allow for the generation of more adversarial examples.
-Specifically, adversarial training defense methods like InfoBERT and FreeLB++ show poor performance on longer texts (more than 300 tokens), with the defense rate reduced to near 0.
-This drastic decline indicates that given an adequate number of queries, the attacker is guaranteed to find a successful attack to fool these models.
-Similarly, EarlyRobust exhibits a performance drop on long texts as it is based on FreeLB training.
-RSMI, however, performs worse on short texts, but its defense rate increases as the text length grows.
-Compared to all SOTA defense approaches, the two DiffuseDef variants show a more steadily declining trend and maintain a higher defense rate across all token lengths, i.e.
-DiffuseDef is more robust to input text length.
+The introduction of randomness in hidden states results in uncertainty in model logits, thus increasing the difficulty of finding vulnerable words to attack (Wang et al., 2023).
+5.2 Robustness against White-box Attacks
+In addition to black-box attacks, we also measure the robustness against white-box attacks, where the attackers can access the defender model’s gradients and parameters, which is a more challenging adversarial attack scenario than black-box attacks.
+In Table 2, we report AUA score under T-PGD (Yuan et al., 2023) and SemAttack Wang et al.
+(2022), the former being a variant of PGD attack (Madry et al., 2019) for text, while the latter adopts C&W attack (Carlini and Wagner, 2017) on embedding layers.
+We observe that white-box attacks result in lower AUA for all defense methods.
+For example, FreeLB++ performs poorly on IMDB when attacked by SemAttack, with the AUA score being 0.03.
+However, applying diffusion training still significantly improves adversarial robustness to white-box attacks.
+Note that even RSMI shows very robust performance against T-PGD444RSMI is incompatible with SemAttack because it requires gradients of the embedding layer, but SemAttack skips the embedding layer by sending embeddings directly to the target model., it can still further improve from DiffuseDef.
+| Method | AGNews | IMDB | ||
+| T-PGD | Sem | T-PGD | Sem | |
+| Fine-Tuned | 8.8 | 41.5 | 3.0 | 1.3 |
+| InfoBERT | 32.1 | 47.1 | 25.4 | 2.0 |
+| FreeLB++ | 19.6 | 58.2 | 15.5 | 3.4 |
+| RSMI | 79.6 | n/a | 43.3 | n/a |
+| DiffuseDef-FreeLB | 59.4 | 68.1 | 50.3 | 28.2 |
+| DiffuseDef-RSMI | 81.7 | n/a | 48.4 | n/a |
+5.3 Ablation analysis
+To understand how each component contributes to DiffuseDef, we conduct an ablation analysis on the AGNews dataset by disabling ensembling, diffusion denoising, and adversarial training one at a time.
+Table 3 shows the ablation results, and we find that removing either component (e.g., ensembling and diffusion denoising) at inference time leads to a sharp decrease in accuracy under attack, which indicates that both denoising and ensembling contribute significantly to the improved robustness.
+Nevertheless, removing adversarial training, though also introducing a degraded AUA score, is less detrimental than the other inference time components, showing that DiffuseDef is not dependent on a specific adversarial training method.
+| Method | TF | TB | BA |
+| DiffuseDef | 84.5 | 86.0 | 84.6 |
+| w/o ensembling | 64.2 | 65.8 | 54.9 |
+| w/o diffusion denoising | 57.1 | 58.4 | 47.7 |
+| w/o adv training | 80.3 | 80.9 | 80.5 |
+5.4 DiffuseDef + Adversarial Augmentation
+In Figure 2, we combine DiffuseDef with adversarial data augmentation and evaluate the robustness on AGNews.
+We first generate adversarial augmentation data using TextFooler and adversarially train the classifier with augmented training data.
+The augmented classifier is further performed with diffusion training, resulting in a DiffuseDef variant.
+We notice that adversarial training on augmented data helps improve the performance on all three attacks.
+After applying diffusion training, the adversarial performance significantly improves over the augmented model, which proves the flexibility of DiffuseDef to be applied with different types of adversarial training methods.
 5.5 Effect of Additional Denoising Steps
-In Figure 3 we study how the inference denoising steps can affect the adversarial performance.
-For the DiffuseDef model without ensembling, both AUA score and the number of queries required to attack increase as the inference denoising step is larger.
+In Figure 3, we study how the inference denoising steps can affect the adversarial performance.
+For the DiffuseDef model without ensembling, both the AUA score and the number of queries required to attack increase as the inference denoising step is larger.
 As the denoising step grows from 1 to 30, the AUA score improves from 58 to 65 while the number of attack queries grows from 430 to 780.
@@
 5.6 Ensembling Diffused Hidden Representations
-In DiffuseDef the text hidden state is diffused and ensembled to form a denoised hidden representation, which contributes significantly to the improved adversarial robustness.
+In DiffuseDef, the text hidden state is diffused and ensembled to form a denoised hidden representation, which contributes significantly to the improved adversarial robustness.
 In this section, we study how the ensembling diffused hidden representation helps defend against adversarial attacks.
 As mentioned in Section 2.1, attack methods need to first rank token importance based on its influence on prediction.
-Specifically, the importance score is calculated by comparing the change of model prediction probablities after removing each word.
-In Figure 4, we compare the distribution of max token importance score between FreeLB++ and its DiffuseDef counterpart.
-Both FreeLB++ and DiffuseDef show a long-tail distribution with over 80 percent examples having a max token importance score below 0.1.
-This suggests that in most cases changing one single token will not significantly alter the prediction for both models.
-However, DiffuseDef shows a notably lower percentage of tokens when the max importance score is between 0.9 and 1, where the attacker can easily find the vulnerable token to construct adversarial examples.
+Specifically, the importance score is calculated by comparing the changes in model prediction probabilities after removing each word.
+In Figure 4, we compare the density distribution of the max token importance score between FreeLB++ and its DiffuseDef counterpart.
+DiffuseDef shows a notably lower percentage than FreeLB++ when the max importance score is high, where the attacker can easily find the vulnerable token to construct adversarial examples.
 This difference shows that DiffuseDef can complicate the process of important word searching, which accounts for the increased number of queries required for a successful attack.
@@
 5.7 Efficiency of DiffuseDef
-| Method | Params | FLOPS |
-| Fine-Tuned (BERT) | 110M | 46G |
-| EarlyRobust | 82M | 32G |
-| FreeLB++ | 110M | 46G |
-| InfoBERT | 110M | 46G |
-| RSMI | 110M | 92G |
-| RanMask () | 110M | 459G |
-| SAFER ( | 110M | 459G |
-| DiffuseDef () | 120M | 96G |
-| DiffuseDef () | 120M | 267G |
+| Method | Params | FLOPS | Train |
+| Fine-Tuned (BERT) | 110M | 46G | 1x |
+| EarlyRobust | 82M | 32G | 0.8x |
+| FreeLB++ | 110M | 46G | 10.5x |
+| InfoBERT | 110M | 46G | 6.5x |
+| RSMI | 110M | 92G | 1.25x |
+| RanMask () | 110M | 459G | 1.2x |
+| SAFER ( | 110M | 459G | 1x |
+| DiffuseDef () | 120M | 96G | 1.1x |
+| DiffuseDef () | 120M | 267G | 1.1x |
 Given that DiffuseDef adds additional denoising and ensembling steps during inference, it inevitably increases the computation time compared to its base model.
-To study its efficiency, we report the number of model parameters and inference FLOPS in Table 4.
-In addition to the defense methods in Table 1, we also compare the efficiency of DiffuseDef with two other SOTA ensembling-based defense methods, i.e.
-RanMask (Zeng et al., 2023) and SAFER Ye et al.
+To study its efficiency, we report the number of model parameters, inference FLOPS, and training time in Table 5.
+In addition to the defense methods in Table 1, we also compare the efficiency of DiffuseDef with two other SOTA ensembling-based defense methods, i.e., RanMask (Zeng et al., 2023) and SAFER Ye et al.
 (2020).
-All SOTA models have the same number of parameters as the fine-tuned BERT model, except EarlyRobust which applies attention head pruning for better efficiency.
-DiffuseDef, with 1 additional diffusion layer, increases the number of parameters from 110M to 120M.
-DiffuseDef requires more inference FLOPS than non ensembling-based baselines such as FreeLB++ and EarlyRobust.
-With and , the FLOPS for DiffuseDef doubles from 46G to 96G, nevertheless, this number is close to RSMI model (92G FLOPS) as it requires gradient information during inference.
-Despite this increase, DiffuseDef is more efficient than ensembling-based methods like RanMask and SAFER which need to go through a full forward pass for all ensembles.
-With the same ensembling number of 10, both RanMask and SAFER require 459G FLOPS, which is 10x the number for BERT baseline.
-In contrast, even with increased to 5, DiffuseDef can be run faster with 267G FLOPS, showing that it can mitigate the efficiency problem from ensembling while maintaining the benefit of improved robustness.
+All SOTA models have the same number of parameters as the fine-tuned BERT model, except EarlyRobust, which applies attention head pruning for better efficiency.
+DiffuseDef requires more inference FLOPS than non-ensembling-based baselines such as FreeLB++ and EarlyRobust, due to the diffusion steps.
+With and , the FLOPS for DiffuseDef doubles from 46G to 96G, nevertheless, this number is close to the RSMI model as it requires gradient information during inference.
+Despite this increase, DiffuseDef is more efficient than ensembling-based methods like RanMask and SAFER, which require a full forward pass for all ensembles.
+With the same ensembling number of 10, both RanMask and SAFER require 459G FLOPS, which is 10x the number for the BERT baseline.
+In contrast, even with increased to 5, DiffuseDef can be run faster with 267G FLOPS, mitigating the efficiency problem from ensembling while maintaining the benefit of improved robustness.
+Regarding training time, although training the diffusion layer requires more epochs, it costs similar time as training the classifier due to the small size of the diffusion layer.
+By contrast, adversarial training methods such as FreeLB++ and InfoBERT require much longer training time due to increased iterations of back-propagation per training batch.
 6 Conclusions
@@
 During inference, the diffusion layer is used to denoise the adversarial hidden states, which are ensembled to construct a robust text representation.
-Our experiments validate the effectiveness and efficiency of DiffuseDef, which significantly outperforms SOTA on three common adversarial attack methods.
+Our experiments validate the effectiveness and efficiency of DiffuseDef, which significantly outperforms SOTA on common black-box and white-box adversarial attack methods, and maintains the performance on clean texts.
 Analysis shows that DiffuseDef makes it difficult to find vulnerable tokens to attack, and also reduces the difference between the hidden representations of clean and adversarial texts.
-7 Limitations
-Scope
-Our experiments focus on defending against three common black-box adversarial attack methods, while whether DiffuseDef improves model robustness against white-box attacks is unclear.
-White-box attacks have access to model parameters and can utilize gradient information to construct adversarial examples more efficiently than black-box attacks.
-Defending against white-box attacks is more challenging, and we consider this as a future direction of DiffuseDef.
-Comparison with additional approaches
-Due to the length limit, we do not compare against all current approaches. However we do compare with the SOTA methods with best adversarial robustness based on our preliminary experiments.
-Efficiency
-Despite the fact that DiffuseDef is more efficient than existing ensembling-based methods, it still requires more model parameters and inference FLOPS than non-ensembling-based models to achieve a better robustness.
-Future directions of this work might involve efforts to reduce the size of diffusion layer and number of ensembles to make DiffuseDef more efficient.
-8 Ethical Considerations
-In this paper we propose a new method DiffuseDef which uses a diffusion layer as a denoiser to provide robust and efficient text representation.
+Limitations
+Despite the fact that DiffuseDef is more efficient than existing ensembling-based methods, it still requires more model parameters and inference FLOPS than non-ensembling-based models to achieve better robustness.
+Future directions of this work might involve efforts to reduce the size of the diffusion layer and the number of ensembles to make DiffuseDef more efficient.
+Ethical Considerations
+In this paper, we propose a new method DiffuseDef, which uses a diffusion layer as a denoiser to provide robust and efficient text representation.
 We demonstrate that the proposed method could significantly improve the robustness of NLP systems to adversarial attacks.
-However, DiffuseDef cannot defend against all adversarial attacks without limitations (e.g.
-number of perturbed words, semantic similarity between original and adversarial examples).
-Potential risks might include creation of new adversarial attacks devised specifically for DiffuseDef.
+However, DiffuseDef cannot defend against all adversarial attacks without limitations (e.g., number of perturbed words, semantic similarity between original and adversarial examples).
+Potential risks might include the creation of new adversarial attacks devised specifically for DiffuseDef.
 Appendix A Data Preparation
@@
 Table 6 presents the number of examples in train/valid/test splits and the average token length for the three datasets used in the experiments.
-For QNLI and AGNews datasets, we randomly split the training set into our train/valid splits, with a ratio of 0.9/0.1, and use its test set as our test split.
+For QNLI and AGNews datasets, we randomly split the training set into our train/valid splits, with a ratio of 0.9/0.1, and use their test sets as our test split.
 For IMDB dataset, we randomly split the dataset into train/valid/test splits with a ratio of 0.8/0.1/0.1.
@@
 The minimum semantic similarity between the clean text and adversarial text is set to 0.84, with the score computed using Universal Sentence Encoder (Cer et al., 2018).
-The maximum number of candidate substitution from attacker is 50, thus the maximum number of queries where is the number of tokens.
-Finally, the maximum percentage of changed tokens is set to 0.3/0.1/0.2 for AGNews, IMDB, and QNLI dataset respectively.
+The maximum number of candidate substitutions from the attacker is 50, thus the maximum number of queries where is the number of tokens.
+Finally, the maximum percentage of changed tokens is set to 0.3/0.1/0.2 for AGNews, IMDB, and QNLI datasets respectively.
 Appendix C Training
@@
 | 10 | 10 | 10 |
-The details on hyper-parameters of diffusion training can be found in Table 7.
+The details on hyperparameters of diffusion training can be found in Table 8.
 All models are trained on a single RTX A6000 GPU.
-The diffusion training of 100 epochs takes 6/4/3 hours on AGNews, IMDB, QNLI datasets respectively.
+The diffusion training of 100 epochs takes 6/4/3 hours on AGNews, IMDB, and QNLI datasets respectively.
 Appendix D License for Scientific Artifacts
@@
 | RoBERTa (Liu et al., 2019) | MIT License |
-Table 8 lists the scientific artifacts including data, codes, and models used in this paper.
-The use of these artifacts in this paper is consistent with their intended use, i.e.
-for scientific research only.
-The data used in the experiment is in English and does not contain personally identifying info or offensive content.
-Appendix E Example of noising and denoising in DiffuseDef
-Adding and removing noise to hidden states are essential features in DiffuseDef which contribute to the improved adversarial robustness.
-To study how adding or removing noise can affect the semantic meaning of the text, we feed the hidden states to the pretrained BERT model with masked language modeling (MLM) head to generate the text output.
-In Table 9, we present the MLM outputs from hidden states added with different steps of noise and the MLM outputs from noise hidden states denoised with same number of steps.
-In the example shown, with more noise added some semantic information can be lost and replaced by symbols or function words like "." or "the".
-In contrast, denoising for the same number of steps help alleviate such information lost.
-For example, the word "IBM" can be recovered from the noise.
-However, in practise it is not possible to assume number of denoising steps therefore in Table 10 we show the MLM outputs of denoised hidden states directly from clean and adversarial texts.
+Table 9 lists the scientific artifacts, including data, codes, and models used in this paper.
+The use of these artifacts in this paper is consistent with their intended use, i.e., for scientific research only.
+The data used in the experiment is in English and does not contain personally identifying information or offensive content.
+Appendix E Robustness with RoBERTa
+In Table 10, we report the black-box attack results on RoBERTa-based models.
+A similar conclusion holds for BERT-based models, and DiffuseDef significantly outperforms other defense methods on all 3 datasets.
+| Dataset | PLM | Method | Clean% | AUA% | #Query | ||||
+| TF | TB | BA | TF | TB | BA | ||||
+| AGNews | RoBERTa-base | Fine-Tuned | 94.9 | 34.1 | 36.9 | 43.6 | 372 | 396 | 410 |
+| InfoBERT | 95.5 | 40.2 | 45.2 | 48.6 | 392 | 421 | 430 | ||
+| FreeLB++ | 95.4 | 57.5 | 62.9 | 55.9 | 444 | 467 | 447 | ||
+| RSMI | 93.1 | 64.2 | 66.4 | 67.4 | 774 | 861 | 808 | ||
+| DiffuseDef-FreeLB++ (Ours) | 95.3 | 85.6 | 87.6 | 85.3 | 880 | 976 | 906 | ||
+| DiffuseDef-RSMI (Ours) | 92.9 | 82.9 | 83.5 | 82.2 | 905 | 925 | 1047 | ||
+| IMDB | RoBERTa-base | Fine-Tuned | 94.6 | 21.3 | 17.9 | 13.6 | 587 | 671 | 493 |
+| InfoBERT | 94.8 | 30.9 | 27.9 | 21.8 | 681 | 760 | 549 | ||
+| FreeLB++ | 95.3 | 46.0 | 42.1 | 33.9 | 829 | 974 | 637 | ||
+| RSMI | 92.7 | 77.9 | 74.3 | 70.6 | 3443 | 4342 | 2619 | ||
+| DiffuseDef-FreeLB++ (Ours) | 95.0 | 86.2 | 85.9 | 86.8 | 3573 | 4663 | 2941 | ||
+| DiffuseDef-RSMI (Ours) | 92.4 | 84.7 | 84.1 | 84.3 | 3673 | 4782 | 3007 | ||
+| QNLI | RoBERTa-base | Fine-Tuned | 92.8 | 26.6 | 22.5 | 20.3 | 204 | 219 | 188 |
+| InfoBERT | 92.5 | 29.1 | 25.8 | 20.3 | 205 | 223 | 189 | ||
+| FreeLB++ | 92.8 | 33.9 | 27.7 | 20 | 227 | 244 | 200 | ||
+| RSMI | 89.3 | 38.8 | 33.2 | 30.0 | 340 | 375 | 343 | ||
+| DiffuseDef-FreeLB++ (Ours) | 92.7 | 64.6 | 64.3 | 61.5 | 473 | 579 | 524 | ||
+| DiffuseDef-RSMI (Ours) | 88.8 | 57.7 | 55.7 | 53.9 | 469 | 578 | 518 | ||
+Appendix F Robustness w.r.t Token Length
+Figure 5 provides a comparison of the defense rate for different models by token length on the IMDB dataset.
+The defense rate is calculated as the percentage of test examples in which TextFooler fails to construct a successful attack.
+All models except RSMI show a consistent trend that the defense rate declines as the texts lengthen.
+This trend can be attributed to the nature of adversarial attacks, as longer texts allow for the generation of more adversarial examples.
+Specifically, adversarial training defense methods like InfoBERT and FreeLB++ show poor performance on longer texts (more than 300 tokens), with the defense rate reduced to nearly 0.
+This drastic decline indicates that, given an adequate number of queries, the attacker is guaranteed to find a successful attack to fool these models.
+Similarly, EarlyRobust exhibits a performance drop on long texts as it is based on FreeLB training.
+RSMI, however, performs worse on short texts, but its defense rate increases as the text length grows.
+Compared to all SOTA defense approaches, the two DiffuseDef variants show a more steadily declining trend and maintain a higher defense rate across all token lengths, i.e., DiffuseDef is more robust to input text length.
+Appendix G Example of noising and denoising in DiffuseDef
+Adding and removing noise to hidden states are essential features in DiffuseDef, which contribute to the improved adversarial robustness.
+To study how adding or removing noise can affect the semantic meaning of the text, we feed the hidden states to the pretrained BERT model with a masked language modeling (MLM) head to generate the text output.
+In Table 11, we present the MLM outputs from hidden states added with different steps of noise and the MLM outputs from noise hidden states denoised with the same number of steps.
+In the example shown, with more noise added some semantic information can be lost and replaced by symbols or function words like “.” or “the”.
+In contrast, denoising for the same number of steps helps alleviate such information loss.
+For example, the word “IBM” can be recovered from the noise.
+However, in practice, it is not possible to assume the number of denoising steps; therefore in Table 12 we show the MLM outputs of denoised hidden states directly from clean and adversarial texts.
 On clean text, we observe that a higher number of denoising steps can result in more abstraction of the texts.
-For example, more words are replaced with "the" in the MLM outputs as grows.
-However, words related to the topic (e.g.
-"Manchester United", "Liverpool") are kept during the denoising process, thus the model can predict correctly.
-Similarly, the trend of abstraction can be also found on adversarial text while we observe that the denoising can help remove the adversarial noise / perturbation and recover the word "united" from "nation", thus resulting its correct prediction on the adversarial text.
+For example, more words are replaced with “the” in the MLM outputs as grows.
+However, words related to the topic (e.g., “Manchester United”, “Liverpool”) are kept during the denoising process, thus the model can predict correctly.
+Similarly, the trend of abstraction can also be found on adversarial text while we observe that the denoising can help remove the adversarial noise/perturbation and recover the word “united” from “nation”, thus resulting in its correct prediction on the adversarial text.
 | MLM Output (add noise) | MLM Output (add noise then denoise) | |
@@
 | Sports | Sports |
-Appendix F Confusion Matrix under Attack
-Figure 5 and 6 present the confusion matrixes of models prediction on clean text and on adversarial texts (successful attack example) on AGNews and IMDB test sets respectively.
+Appendix H Confusion Matrix under Attack
+Figure 6 and 7 present the confusion matrices of models’ predictions on clean text and on adversarial texts (successful attack example) on AGNews and IMDB test sets respectively.